# MambaCSR: Dual-Interleaved Scanning for Compressed Image Super-Resolution With SSMs

### v1

@@ 2.2 State Space Model in Vision Tasks
Recently, the State Space Model has demonstrated high efficiency in capturing the dynamics and dependencies of long sequences. This has garnered significant attention in the visual field, covering various domains including 3D visual recognition [67, 27], medical imaging [33, 54], multi-modal understanding [55, 58], remote sensing images [59, 16], and more. Due to its impressive potential performance gains, SSM has also been introduced into low-level vision tasks [15, 21, 73, 50]. Among these, MambaIR [15] employs 2D Selective Scanning (SS2D) [30] as the backbone while incorporating channel attention and convolutional layers to enhance local information capabilities. In contrast, VMambaIR [42], is based on a U-Net structure and extends channel dimension scanning for further feature extraction. FreqMamba [70] operates from the frequency domain, utilizing Mamba blocks to achieve improved de-raining effects. LFMamba [31] employs an efficient SS2D mechanism to achieve light-field image-super-resolution. However, most existing Mamba approaches [7] in the low-level field are preliminary explorations, relying primarily on the sequential scanning method. This approach overlooks the control over local information, which is critical in low-level tasks. In this paper, we introduce a dual-interleaved scanning method and a cross-scale scanning method to enhance the impact of local pixels and multi-scale features.
@@ 3.1 Preliminaries
Finally, the output could be obtained through a global convolution as follows: ($\bar{K}$ denotes the convolution kernel.)
$y=x\odot\bar{K},\quad\bar{K}=\left(\bar{C}\bar{B},\,\bar{C}\bar{A}\bar{B},\,\ldots,\,\bar{C}\bar{A}^{L-1}\bar{B}\right).$ (3)
@@ 3.2 Motivations
As shown in Fig. 2(b), we employ a dual-interleaved scanning approach to enhance the modeling of local dependencies and long-range contextual information, while also improving computational efficiency. This approach interleaves two types of scanning: (i) Hierarchical interleaved scanning is applied every two blocks to balance local and global feature extraction. (ii) Interleaved horizontal-to-vertical scanning. The original scanning method includes four scans, which introduces computational redundancy; consequently, one scanning path might dominate the critical information, while the others contribute less (Shi, Dong, and Xu 2024). To address this, we reduce the four directional scans to two horizontal and two vertical scans, with each block undergoing only two scans. Crucially, we ensure that each scan is paired with its flipped counterpart. This pairing is necessary due to the non-causality of Mamba, where only preceding tokens contribute to subsequent tokens within a sequence. Paired scanning is therefore essential to ensure bidirectional information flow.
@@ 3.5 Network Architecture
The Residual Local Mamba Block (RLMB) focuses on deep modeling of input features, capturing both local and long-range information. In each block, we first apply a normalization layer, followed by the adoption of a local-window-based 2D Selective Scan Module (LW-SS2D) to capture spatial long-range dependencies, as shown in Fig. 1 (b). We use different window sizes to represent local window-based and sequential scanning methods. For an input image size of $64\times 64$ during the training phase, the window size for local scanning is set to 8, while the window size for sequential scanning is set to 64. Inspired by work HAT [5], we also employ channel attention block parallel to LW-SS2D to further enhance the representation ability of the network. Followed by MambaIR [15], we use a scaling factor $s_{1}$ to control the intensity of input feature $X$ through skip connection. After establishing long-term relationships, we apply Layer normalization and MLP to process the intermediate features. Scaling factor $s_{2}$ is also used to control the magnitude of residual. Thus, the entire process of RLMB could be expressed as:
@@
Methods Params (M) GFLOPS PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
RCAN 15.59 65.25 24.87 0.7019 23.72 0.5983 22.65 0.7128 21.58 0.5795 24.77 0.6639
SwinIR 8.72 37.17 25.10 0.7152 23.97 0.6027 23.31 0.7354 22.03 0.6004 25.61 0.6861
Swin2SR 11.67 52.96 25.12 0.7162 23.96 0.6028 23.35 0.7358 22.08 0.6008 25.64 0.6868
HST 16.48 54.95 25.12 0.7156 24.00 0.6034 23.39 0.7365 22.10 0.6026 25.66 0.6887
HAT 20.81 102.4 25.13 0.7177 24.02 0.6052 23.42 0.7362 22.13 0.6037 25.66 0.6895
MambaIR 16.70 82.30 25.13 0.7176 24.01 0.6054 23.46 0.7379 22.14 0.6041 25.68 0.6894
MambaCSR (w/o) 16.91 75.01 25.16 0.7201 24.06 0.6065 23.52 0.7400 22.22 0.6071 25.72 0.6906
MambCSR (Final) 19.18 80.46 25.20 0.7209 24.07 0.6072 23.57 0.7419 22.26 0.6081 25.74 0.6911
[…]
4 Experiments
@@ 4.1 Dataset and Implementation Details
We use the DF2K [1] dataset as our primary training dataset. Given that JPEG [46] is currently the most widely used compression standard, we apply JPEG compression with quality factors from range [10, 20, 30] to the dataset using the JPEG codec function provided by the OpenCV library. For evaluation, we test on five commonly used super-resolution benchmarks: Set5 [2], Set14 [63], Manga109 [34], Urban100 [18], and the DIV2K test set [1]. The entire implementation is based on the PyTorch framework. During training, we employ the Adam optimizer ($\beta_{1}$= 0.9, $\beta_{2}$ = 0.99) with an initial learning rate of 2e-4, which is decayed by a factor of 0.5 at 150k, 225k, and 275k iterations. Due to the small size of the DF2K dataset and its rapid convergence, the total number of iterations is set to 300k. Input low-resolution images are randomly cropped to 64 $\times$ 64 and augmented with random flips and rotations. The window size for interleaved scanning is set to 8. The total batch size is set to 32, trained on eight NVIDIA V100 GPUs.
@@ 4.2 Comparison with State-of-the-Art
We compare our proposed MambaCSR with three image super-resolution models: RCAN [68], SwinIR [28], and HAT [5]. Additionally, we include two state-of-the-art compressed image super-resolution models: Swin2SR [8] and HST [23], as well as the SSM-based baseline model, MambaIR [15]. All models are trained on the compressed DF2K dataset using the same settings. For evaluation, we employ PSNR and SSIM metrics, with PSNR values specifically measured on the Y channel.
Figure 3: Visual comparison of our MambaCSR with other SOTA models on Urban100 with scale x4.
Quantitative Results. From Table 1, it is evident that our proposed MambaCSR achieves the best performance across all five compressed benchmarks in terms of PSNR and SSIM. Specifically, on the Manga109 and Urban100 datasets, MambaCSR with cross-scale scanning outperforms MambaIR by an average of 0.12 dB, highlighting the effectiveness of our framework. We attribute this improvement to the introduction of dual-interleaved scanning, which enhances the model’s focus on local information, and the use of cross-scale scanning, which better leverages multi-scale features for removing compression and downsampled artifacts. Additional results under different quality factor settings [20, 30] are provided in the Sec. A.3 of supplementary.
Qualitative Results. We present visual comparisons with other models. As shown in Fig. 3, our proposed MambaCSR excels in handling compression artifacts, particularly in restoring textures and fine details, highlighting the robustness of our model.
Computational Comparisons. To demonstrate the efficiency of our dual-interleaved scanning method, we evaluate the parameter count and computational complexity of the proposed modules and compare them with other models. As shown in Table 1, the computational cost of using only the dual-interleaved scanning method is 10 GFLOPS lower than that of MambaIR. Even when incorporating cross-scale scanning, our computational complexity remains comparable to MambaIR, highlighting the efficiency of the dual-interleaved scanning method.
Figure 4: The receptive fields of different scanning methods: (a) Sequential scanning adopted in MambaIR, (b) Our hierarchical interleaved scanning, (c) Only local window-based scanning. All results are tested on the Manga109 dataset.
@@ 4.3 Ablation Studies
  Manga109 Urban100
Scanning Methods PSNR SSIM PSNR SSIM
No scanning 23.53 0.7400 22.22 0.6071
Position-Aligned 23.57 0.7419 22.26 0.6081
Sequential Scanning 23.53 0.7403 22.24 0.6077
Table 3: Effects of two cross-scale scanning methods: (I) our position-aligned cross-scale scanning. (II) simple sequential scanning.
Effects of different cross-scale scanning methods. In this paper, we introduce the position-aligned cross-scale scanning method, as shown in Fig. 2(c). To validate the effectiveness of our position-aligned approach, we also implement a sequential scanning method, where all tokens of the low-resolution image are scanned first, followed by scanning the tokens of the high-resolution image. We then compare the performance of these two approaches. As shown in Table 3, the position-aligned cross-scale scanning method outperforms the sequential scanning approach. This indicates that the position-aligned method is more effective in modeling multi-scale features. The likely reason for the underperformance of the simple sequential scanning approach is that the long-sequence decay properties inherent to the SSM reduce the influence of low-resolution tokens on the original feature tokens.
Effects of dual-interleaved scanning and CAB. To validate the effectiveness of our proposed dual-interleaved scanning method and the channel attention block used in each RLMB, we separately test the hierarchical interleaved scanning and horizontal-to-vertical interleaved scanning components within the dual-interleaved framework. As shown in Table 4, hierarchical interleaved scanning yields the highest performance gain, followed by the channel attention block. Although the performance of horizontal-to-vertical scanning alone shows a slight decrease, its integration with the other two methods maintains comparable performance while offering the advantage of reduced computational complexity. This demonstrates the capability of the dual-interleaved scanning method to balance performance gains with reduced computational complexity. We also present the receptive fields of different scanning methods, including sequential scanning, window-based scanning, and our proposed hierarchical-interleaved scanning. As depicted in Fig. 4, the sequential scanning method exhibits a smaller central receptive field while maintaining some level of global information perception. In contrast, window-based scanning enhances the local receptive field in the center but loses some capacity for global information modeling. Our hierarchical interleaved scanning method, on the other hand, balances local and global receptive fields, retaining the long-range modeling capacity while providing a larger local receptive field compared to the sequential scanning method.
  Hierarchical Horizontal-to-vertical CAB Manga109 Urban100
23.44 22.12
23.49 22.18
23.43 22.11
23.47 22.15
23.48 22.16
23.53 22.22
23.47 22.14
23.52 22.20
Table 4: Effects of hierarchical interleaved scanning, horizontal-to-vertical interleaved scanning and CAB. The performance is tested on Manga109. The best performance is in bold.
@@ A.1 Broader Impact
MambaCSR, as proposed in this paper, introduces an innovative dual-interleaved scanning paradigm for compressed image super-resolution (CSR). By leveraging hierarchical interleaved scanning and horizontal-to-vertical scanning, MambaCSR significantly reduces redundancy while enhancing contextual information modeling, leading to more efficient and accurate super-resolution. These advancements not only improve the performance of CSR tasks in industrial applications but also contribute to reducing the computational burden associated with processing high-resolution images. The primary applications of MambaCSR span a range of areas, including image restoration, multimedia content enhancement, and visual communication. Importantly, the deployment of MambaCSR adheres to ethical standards, ensuring that the technology is used responsibly to avoid potential biases and harms in visual data processing. By focusing on efficiency and accuracy, MambaCSR paves the way for more sustainable and equitable advancements in the field of image processing.

### v2

@@ 2.2 State Space Model in Vision Tasks
Recently, the State Space Model has demonstrated high efficiency in capturing the dynamics and dependencies of long sequences. This has garnered significant attention in the visual field, covering various domains including 3D visual recognition [73, 32], medical imaging [38, 60], multi-modal understanding [61, 64], remote sensing images [65, 22], and more. Due to its impressive potential performance gains, SSM has also been introduced into low-level vision tasks [21, 27, 79, 56, 15, 11, 14, 1, 20]. Among these, MambaIR [21] employs 2D Selective Scanning (SS2D) [35] as the backbone while incorporating channel attention and convolutional layers to enhance local information capabilities. In contrast, VMambaIR [47], is based on a U-Net structure and extends channel dimension scanning for further feature extraction. FreqMamba [76] operates from the frequency domain, utilizing Mamba blocks to achieve improved de-raining effects. LFMamba [36] employs an efficient SS2D mechanism to achieve light-field image-super-resolution. However, most existing Mamba approaches [8] in the low-level field are preliminary explorations, relying primarily on the sequential scanning method. This approach overlooks the control over local information, which is critical in low-level tasks. In this paper, we introduce a dual-interleaved scanning method and a cross-scale scanning method to enhance the impact of local pixels and multi-scale features.
@@ 3.1 Preliminaries
Mamba aims to dynamically select $\bar{A}$ and $\bar{B}$ based on the input, incorporating parallel scanning to enhance computational efficiency.
@@ 3.2 Motivations
As shown in Fig. 2(b), we employ a dual-interleaved scanning approach to enhance the modeling of local dependencies and long-range contextual information, while also improving computational efficiency. This approach interleaves two types of scanning: (i) Hierarchical interleaved scanning is applied every two blocks to balance local and global feature extraction. (ii) Interleaved horizontal-to-vertical scanning. The original scanning method includes four scans, which introduces computational redundancy; consequently, one scanning path might dominate the critical information, while the others contribute less (The reason could be found in Supplementary). To address this, we reduce the four directional scans to two horizontal and two vertical scans, with each block undergoing only two scans. Crucially, we ensure that each scan is paired with its flipped counterpart. This pairing is necessary due to the non-causality of Mamba, where only preceding tokens contribute to subsequent tokens within a sequence. Paired scanning is therefore essential to ensure bidirectional information flow.
@@ 3.5 Network Architecture
The Residual Local Mamba Block (RLMB) focuses on deep modeling of input features, capturing both local and long-range information. In each block, we first apply a normalization layer, followed by the adoption of a local-window-based 2D Selective Scan Module (LW-SS2D) to capture spatial long-range dependencies, as shown in Fig. 1 (a). We use different window sizes to represent local window-based and sequential scanning methods. For an input image size of $64\times 64$ during the training phase, the window size for local scanning is set to 8, while the window size for sequential scanning is set to 64. Inspired by work HAT [6], we also employ channel attention block parallel to LW-SS2D to further enhance the representation ability of the network. Followed by MambaIR [21], we use a scaling factor $s_{1}$ to control the intensity of input feature $X$ through skip connection. After establishing long-term relationships, we apply Layer normalization and MLP to process the intermediate features. Scaling factor $s_{2}$ is also used to control the magnitude of residual. Thus, the entire process of RLMB could be expressed as:
@@ 4 Experiments
Codec Methods Params GFLOPS PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SwinIR 8.72 37.17 25.10 0.715 23.97 0.603 23.32 0.735 22.03 0.600 25.61 0.686
Swin2SR 11.67 52.96 25.12 0.716 23.96 0.603 23.34 0.736 22.05 0.601 25.64 0.687
HST 16.48 54.95 25.12 0.716 24.00 0.603 23.35 0.737 22.09 0.602 25.65 0.688
HAT 20.81 102.4 25.13 0.718 24.02 0.604 23.42 0.784 22.13 0.603 25.66 0.690
MambaIR 16.70 82.3 25.14 0.718 24.01 0.605 23.46 0.738 22.14 0.604 25.68 0.689
JPEG (QF=10) MambaCSR 19.69 76.03 25.20 0.721 24.07 0.607 23.57 0.742 22.27 0.608 25.74 0.691
SwinIR 8.72 37.17 26.51 0.762 25.03 0.639 24.89 0.780 23.04 0.646 26.72 0.718
Swin2SR 11.67 52.96 26.56 0.763 25.05 0.639 24.92 0.781 23.07 0.647 26.73 0.718
HST 16.48 54.95 26.53 0.762 25.02 0.639 24.91 0.780 23.06 0.646 26.73 0.718
HAT 20.81 102.4 26.49 0.762 25.04 0.639 24.98 0.784 23.12 0.649 26.74 0.719
MambaIR 16.70 82.3 26.61 0.764 25.11 0.640 24.99 0.783 23.13 0.653 26.76 0.720
JPEG (QF=20) MambaCSR 19.69 76.03 26.67 0.768 25.13 0.642 25.12 0.788 23.24 0.658 26.79 0.720
SwinIR 8.72 37.17 27.48 0.784 25.61 0.659 25.79 0.805 23.59 0.672 27.28 0.735
Swin2SR 11.67 52.96 27.51 0.785 25.61 0.659 25.81 0.806 23.63 0.673 27.30 0.735
HST 16.48 54.95 27.49 0.785 25.60 0.659 25.8 0.805 23.61 0.676 27.31 0.736
HAT 20.81 102.4 27.53 0.785 25.61 0.660 25.84 0.807 23.67 0.677 27.32 0.736
MambaIR 16.70 82.3 27.56 0.786 25.64 0.661 25.87 0.808 23.72 0.679 27.34 0.737
JPEG (QF=30) MambaCSR 19.69 76.03 27.63 0.788 25.70 0.663 26.01 0.812 23.82 0.6825 27.39 0.739
SwinIR 8.72 37.17 29.04 0.819 26.71 0.705 27.99 0.856 24.76 0.723 28.96 0.799
Swin2SR 11.67 52.96 29.03 0.818 26.68 0.703 28.01 0.855 24.79 0.725 28.94 0.798
HST 16.48 54.96 29.06 0.819 26.71 0.703 27.97 0.853 24.82 0.726 28.95 0.798
HAT 20.81 102.4 29.11 0.821 26.78 0.706 28.15 0.861 24.87 0.731 28.98 0.801
MambaIR 16.70 82.3 29.12 0.822 26.81 0.707 28.17 0.861 24.89 0.732 29.00 0.802
HEVC (QP=32) MambaCSR 19.69 76.03 29.21 0.824 26.86 0.709 28.28 0.867 25.02 0.737 29.04 0.803
SwinIR 8.72 37.17 27.12 0.774 25.41 0.653 23.55 0.673 26.11 0.814 26.84 0.730
Swin2SR 11.67 52.96 27.13 0.773 25.42 0.652 23.58 0.672 26.12 0.813 26.83 0.728
HST 16.48 54.96 27.09 0.773 25.39 0.651 23.54 0.671 26.09 0.811 26.83 0.728
HAT 20.81 102.4 27.19 0.776 25.46 0.655 23.72 0.678 26.21 0.819 26.87 0.601
MambaIR 16.70 82.3 27.18 0.777 25.46 0.655 23.74 0.680 26.22 0.818 26.89 0.732
HEVC (QP=37) MambaCSR 19.69 76.03 27.22 0.779 25.51 0.658 23.83 0.682 26.35 0.822 26.94 0.735
SwinIR 8.72 37.17 24.82 0.713 23.80 0.598 22.10 0.607 23.81 0.752 23.55 0.551
Swin2SR 11.67 52.96 24.79 0.709 23.78 0.597 22.11 0.607 23.79 0.753 23.52 0.551
HST 16.48 54.96 24.80 0.711 23.77 0.597 22.09 0.606 23.78 0.750 23.55 0.550
HAT 20.81 102.4 24.86 0.714 23.83 0.601 22.24 0.614 23.92 0.757 23.59 0.553
MambaIR 16.70 82.3 24.88 0.714 23.83 0.731 22.22 0.613 23.94 0.758 23.61 0.554
HEVC (QP=42) MambaCSR 19.69 76.03 24.93 0.719 23.87 0.604 22.33 0.618 24.03 0.762 23.64 0.554
@@ 4.1 Dataset and Implementation Details
We use the DF2K [2] dataset as our primary training dataset. The DF2K training set is compressed using three widely used codecs: JPEG [51], HEVC [48], and VVC [4]. For each codec, we select three bitrate points as follows: (i) JPEG: quality factors (QF) of $[10,20,30]$, using the JPEG function provided by the OpenCV library; (ii) HEVC: quantization parameters (QP) of $[32,37,42]$, with compression conducted via the HM software; and (iii) VVC: QP values of $[32,37,42]$, using the VTM software for compression. For evaluation, we test on five commonly used super-resolution benchmarks: Set5 [3], Set14 [69], Manga109 [39], Urban100 [24], and the DIV2K test set [2]. The entire implementation is based on the PyTorch framework. During training, we employ the Adam optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.99$) with an initial learning rate of $2e^{-4}$, which is decayed by a factor of 0.5 at $150k$, $225k$, and $275k$ iterations. Due to the small size of the DF2K dataset and its rapid convergence, the total number of iterations is set to $300k$. Input low-resolution images are randomly cropped to 64 $\times$ 64 and augmented with random flips and rotations. The window size for interleaved scanning is set to 8. The total batch size is set to 32, trained on eight NVIDIA V100 GPUs.
@@ 4.2 Comparison with State-of-the-Art
We compare our proposed MambaCSR with two image super-resolution models (SwinIR [33] and HAT [6]), two compressed image super-resolution models (Swin2SR [9] and HST [29]), and the Mamba-based baseline model, MambaIR [21]. All models are trained on the compressed DF2K dataset under the same settings. For evaluation, we use PSNR and SSIM metrics, with PSNR measured specifically on the Y channel.
Figure 3: Visual comparison of our MambaCSR model with other state-of-the-art models on the Urban100 [24] dataset under the JPEG codec at compression levels QF = [10, 20, 30]. Additional visual comparisons for HEVC and VVC codecs could be found in Supplementary.
Quantitative Results. From Table 1, it is clear that our proposed MambaCSR achieves the highest performance across all five compressed benchmarks and at all compression levels for each codec in terms of PSNR and SSIM. Notably, on the Manga109 and Urban100 datasets, MambaCSR with cross-scale scanning surpasses MambaIR by an average of 0.12 dB, underscoring the effectiveness of our framework. This improvement is attributed to the introduction of dual-interleaved scanning, which enhances the model’s focus on local information, and cross-scale scanning, which better leverages multi-scale features to mitigate compression and downsampling artifacts. Additional results for VVC codec compression with QP values of [32, 37, 42] are provided in the Supplementary.
Qualitative Results. We present visual comparisons with other models, showcasing three different compression levels (QF = [10, 20, 30]) under the JPEG codec. Additional visual comparisons for the HEVC and VVC codecs could be found in the Supplementary. As shown in Fig. 3, our proposed MambaCSR demonstrates superior performance in handling compression artifacts, particularly in restoring textures and fine details, underscoring the robustness of our model.
Computational Comparisons. As shown in Table 1, the computational cost of our proposed MambaCSR is 6 GFLOPS lower than that of MambaIR. Additionally, as shown in Table 2, when cross-scale scanning is omitted and only dual-interleaved scanning is applied, GFLOPS are reduced by approximately 11.
Figure 4: The receptive fields of different scanning methods: (a) Sequential scanning adopted in MambaIR, (b) Our hierarchical interleaved scanning, (c) Only local window-based scanning. All results are tested on the Manga109 [39] dataset.
@@ 4.3 Ablation Studies
Table 2: Effects of two cross-scale scanning methods: (I) our position-aligned cross-scale scanning. (II) simple sequential scanning. No scanning means only using dual-interleaved scanning.
  Scanning Methods GFLOPs Manga109 Urban100
PSNR SSIM PSNR SSIM
No scanning 70.45 23.53 0.7400 22.22 0.6071
Position-Aligned 76.03 23.57 0.7419 22.26 0.6081
Simple scanning 76.03 23.53 0.7403 22.24 0.6077
Effects of different cross-scale scanning methods. In this paper, we introduce the position-aligned cross-scale scanning method, as shown in Fig. 2(c). To validate the effectiveness of our position-aligned approach, we also implement a simple sequential scanning method, where all tokens of the low-resolution image are scanned first, followed by scanning the tokens of the high-resolution image. We then compare the performance of these two approaches. As shown in Table 2, the position-aligned cross-scale scanning method outperforms the sequential scanning approach. This indicates that the position-aligned method is more effective in modeling multi-scale features. The likely reason for the underperformance of the simple sequential scanning approach is that the long-sequence decay properties inherent to the SSM reduce the influence of low-resolution tokens on the original feature tokens.
Effects of dual-interleaved scanning. To validate the effectiveness of our dual-interleaved scanning method, we separately test the hierarchical interleaved scanning and horizontal-to-vertical interleaved scanning components within this framework. As shown in Table 3, hierarchical interleaved scanning yields the greatest performance gain, while horizontal-to-vertical scanning reduces 8 GFLOPs while maintaining comparable performance. The combination of these methods achieves a balance between performance and computational efficiency, demonstrating the dual-interleaved scanning method’s ability to enhance performance with reduced complexity. Additionally, we examine the receptive fields of different scanning methods, including sequential scanning, window-based scanning, and our proposed hierarchical-interleaved scanning. As illustrated in Fig. 4, sequential scanning offers a smaller central receptive field while preserving some global information. Window-based scanning improves the central local receptive field but sacrifices global information modeling. In contrast, our hierarchical interleaved scanning effectively balances local and global receptive fields, retaining long-range modeling capacity while providing a larger local receptive field compared to sequential scanning.
Table 3: Effects of hierarchical interleaved scanning and horizontal-to-vertical interleaved scanning. The performance is tested on Manga109 and Urban100 datasets. The best performance is in bold.
  Hierarchical Horizontal-to-vertical GFLOPs Manga109 Urban100
84.56 23.44 22.12
84.56 23.52 22.20
76.03 23.43 22.12
76.03 23.53 22.22
@@ A.1 Broader Impact
We propose MambaCSR in this paper, which introduces an innovative dual-interleaved scanning paradigm for compressed image super-resolution (CSR). By leveraging hierarchical interleaved scanning and horizontal-to-vertical scanning, MambaCSR significantly reduces redundancy while enhancing contextual information modeling, leading to more efficient and accurate super-resolution. These advancements not only improve the performance of CSR tasks in industrial applications but also contribute to reducing the computational burden associated with processing high-resolution images. The primary applications of MambaCSR span a range of areas, including image restoration, multimedia content enhancement, and visual communication. Importantly, the deployment of MambaCSR adheres to ethical standards, ensuring that the technology is used responsibly to avoid potential biases and harms in visual data processing. By focusing on efficiency and accuracy, MambaCSR paves the way for more sustainable and equitable advancements in the field of image processing.